Ask
your own question, for FREE!
Geometry
7 Online
A fair coin is tossed 3 times in a row. What is the probability that heads appears on only the last toss?
Still Need Help?
Join the QuestionCove community and study together with friends!
@Kitten_is_back
@YoungStudier
#medal given
okay
soooo..... what is the answer?
Still Need Help?
Join the QuestionCove community and study together with friends!
holf on
@ajo21
@Error1603
1/8 is the answer
Your right but how did you get it?
Still Need Help?
Join the QuestionCove community and study together with friends!
P(2 heads) = 3/8 = 1/4 = 0.375 = 37.5% (HHT, HTH, THH) P(at least 2 heads) = P(2 heads) + P(3 heads) = 3/8 + 1/8 = 4/8 = 0.5 = 50% P(at least 1 head) = P(1H) + P(2H) + P(3H) = 3/8 + 3/8 + 1/8 = 7/8 = 0.875 = 87.5% P(1 head) = 3/8 = 0.375 = 37.5% (HTT, THT, TTH) P(1 head on last toss) = 1/8 = 0.125 = 12.5% (TTH) @klc097
just poped up in m u hwad
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
21 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal